OpenStudy (anonymous):

the ratio of red candies to green in a bag is 4/5.which ratio would form a proportional relationship with this ratio

OpenStudy (anonymous):

isnt that 1:4

OpenStudy (anonymous):

The options are 16/25 20/25 20/35 & 24/32

OpenStudy (anonymous):

@kantalope

OpenStudy (anonymous):

oh so the fraction just needs to be equivalent 4/5 != 16/25 4x4=16 ; but 4x5 != 25 but it looks like the next one would work

OpenStudy (anonymous):

Thanks so much! i understand a little more

OpenStudy (mathstudent55):

Multiply both the numerator and denominator of the given ratio by 2, then by 3, then by 4, etc. You get this: 4/5 = 8/10 = 12/15 = 16/20 = 20/25 = 24/30 That means all these fractions are in the same ratio as 4/5. Pick the one you see in your choices.